OpenStudy (sillybutt100):

I need help with simplifying rational expressions. The problem will be in the reply section of the question. I can get most of the problem done but i don't know how to combine the two parts.

OpenStudy (sillybutt100):

\[\frac{ 1 }{ \sqrt{6} } +\frac{ 3 }{ \sqrt{96} }\]

OpenStudy (sillybutt100):

I am able to get it to \[\frac{ 1\sqrt{6} }{ 6 } + \frac{ 1\sqrt{96} }{ 32 }\] but im not sure how to combine them into on exression

OpenStudy (studygurl14):

First off, that is incorrect

OpenStudy (sillybutt100):

what do you mean

OpenStudy (studygurl14):

\[\frac{1}{\sqrt{6}}+\frac{3}{\sqrt{96}}\] \[\frac{1}{\sqrt{6}}+\frac{3}{\sqrt{16\times6}}\] \[\frac{1}{\sqrt{6}}+\frac{3}{4\sqrt{6}}\]

OpenStudy (studygurl14):

\[\frac{4}{4\sqrt{6}}+\frac{3}{4\sqrt{6}}\] Now you can add

OpenStudy (sillybutt100):

\[\frac{ 7 }{ 8\sqrt{6} }\]

OpenStudy (studygurl14):

nope. \[\frac{7}{4\sqrt{6}}\]

OpenStudy (studygurl14):

Now, rationalize the denominator \[\frac{7}{4\sqrt{6}}\times \frac{4\sqrt{6}}{4\sqrt{6}}\] \[\frac{7(4\sqrt{6})}{16(6)}\]

OpenStudy (studygurl14):

\[\frac{28\sqrt{6}}{96}\]

OpenStudy (studygurl14):

\[\frac{7\sqrt{6}}{24}\]

OpenStudy (studygurl14):

the end.

OpenStudy (sillybutt100):

Thank you and while you are here I need to know how to simplify two more kinds of problems. Is that okay?

OpenStudy (studygurl14):

Sure. Please post in a different thread since this one is kinda long :)

OpenStudy (sillybutt100):

ok